OpenStudy (anonymous):

Which ratio forms a proportion with 18/30 ? 20/35 50/90 36/48 60/100

OpenStudy (leader):

18/30=9/15=3/5 if we multiply it by 20/20 we get 60/100

OpenStudy (leader):

So D. :)

OpenStudy (anonymous):

Thank you @Leader I have one more, would you mind helping me with it?

OpenStudy (leader):

Sure post another question :)

OpenStudy (anonymous):

Which proportion has cross products of 5 x 24 and 8 x 15 ? 5/8 = 15/24 5/24 = 8/15 8/24 = 15/5 15/8 = 5/24

OpenStudy (leader):

5/8 = 15/24

OpenStudy (leader):

So A.

OpenStudy (anonymous):

Thank you! :)

OpenStudy (leader):

No prob.